# Predicting Steady-State Behavior in Complex Networks with Graph Neural Networks

### v2

@@ I Introduction
Figure 1: Steady-state behavior in linear dynamical systems on complex networks. We depict the nodes in the graphs with $x-y$ coordinates. We assign the sizes of a node based on the degree of a node. The $z-axis$ portrays the amount of information ($x_{i}^{*}$) on a node in the steady state. The steady-state behavior of linear dynamics on the ER random network leads to delocalization, the Scalefree network shows weak localization, and the star network shows a strong localization.
@@ II Problem Definition
Figure 2: The architecture of the Graph Neural Networks for the regression task over the graphs. The $i^{th}$ input graph (${\bf A}^{(i)}$) and the associated node features (${\bf H}^{(i)}$) are given in matrix form to the models. The Graph Neural Network (GNN) layers of the model output updated node feature matrix (${\bf H}^{{}^{\prime}(i)}$), and the readout layer gives graph level representation as $\bm{z}^{(i)}$. Further, a fully connected layer predicts the IPR value. Finally, we apply a threshold scheme (Eq. (10)) to identify different linear dynamical states.
We quantify the (de)localization behavior of the steady-state ($\bm{x}^{*}$) or the PEV ($\bm{u}\equiv\bm{u}_{1}^{\bf A}$) using the inverse participation ratio ($y_{\bm{x}^{*}}$), which is the sum of the fourth power of the state vector entries and calculate as [20]
@@
where $x^{*}_{i}$ is the $i^{\text{th}}$ component of $\bm{x^{*}}=(x^{*}_{1},x^{*}_{2},\dots,x^{*}_{n})^{T}$ and $\sum_{i=1}^{n}{x^{*}_{i}}^{2}$ is the normalization term. A vector with component $(c,c,\ldots,c)$ is delocalized and has $y_{\bm{x}^{*}}=\frac{1}{n}$ for some positive constant $c>0$, whereas the vector with components $(1,0,\ldots,0)$ yields $y_{\bm{x}^{*}}=1$ and referred as most localized. Furthermore, we consider the networks to be simple, connected and undirected. Hence, some information can easily propagate from one node to another and we never get a steady-state vector of the form $\bm{x}^{*}=(1,0,\ldots,0)$ for a connected network, and thus the IPR value lies between $\frac{1}{n}\leq y_{\bm{x}^{*}}<1$. Therefore, the localization-delocalization behavior of the linear dynamics in the network is quantified using a real value, i.e., each graph associates an IPR value [21, 20, 8, 11]. Now, to identify the states ($\bm{x}^{*}$) belong to which category of dynamical behavior for linear dynamics, we formalize a threshold scheme for identifying IPR values ($y\equiv y_{\bm{x}^{*}}$) lies in the range $[1/n,1)$. We define two thresholds, $\tau_{1}$ and $\tau_{2}$, such that $11/n\leq\tau_{1}<\tau_{2}<$ 1. An additional parameter $\epsilon$ ($\epsilon>0$) defines some flexibility around the thresholds.
@@
$r(y,\tau_{1},\tau_{2},\epsilon)=\begin{cases}1&\text{if }y\leq\tau_{1}-\epsilon\\ 2&\text{if }\tau_{1}-\epsilon<y<\tau_{2}+\epsilon\\ 3&\text{if }y\geq\tau_{2}+\epsilon\\ \end{cases}$ (10)
For instance, we consider a set of threshold values as $\tau_{1}=0.05$, $\tau_{2}=0.2$, $\epsilon=1e-6$. Now, if we consider a regular network (each node have the same degree) of $n$ nodes, we have PEV, $\bm{u}^{\mathcal{R}}=(\frac{1}{\sqrt{n}},\frac{1}{\sqrt{n}},\ldots,\frac{1}{\sqrt{n}})$ of A (Theorem 6 [22]) yielding, $y_{\bm{u}^{\mathcal{R}}}=\frac{1}{n}$, thus $y_{\bm{u}^{\mathcal{R}}}\rightarrow 0$ as $n\rightarrow\infty$. On the other hand, for a star graph having $n$ nodes, $\bm{u}^{\mathcal{S}}=\biggl{(}\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2(n-1)}},\ldots,\frac{1}{\sqrt{2(n-1)}}\biggr{)}$ and, $y_{\bm{u}^{\mathcal{S}}}=\frac{1}{4}+\frac{1}{4(n-1)}$. Hence, for $n\rightarrow\infty$, we get $y_{\bm{u}^{\mathcal{S}}}\approx 0.25$, and PEV is strongly localized for the star networks. Further, it is also difficult to find a closed functional form of PEV for any network, and thereby, it is hard to find the IPR value analytically. For instance, in Erdös-Rényi (ER) random networks, we get a delocalized PEV due to each node having the same expected degree [23]. In contrast, the presence of power-law degree distribution for SF networks leads to some localization in the PEV. For SF networks, the IPR value, while being larger than the ER random networks, is much lesser than the star networks [24]. It may seem that when the network structure is close to regular, linear dynamics are delocalized, and increasing degree heterogeneity increases the localization. However, always looking at the degree heterogeneity not able to decide localization and analyzing structural and spectral properties is essential [20, 11]. A fundamental question at the core of the structural-dynamical relation is: Can we predict the steady state behavior of a linear dynamical process in complex networks?
@@
[…]
Here, we formulate the problem as a graph regression task to predict a target value, IPR, associated with each graph structure. For a given set of graphs $\{\mathcal{G}_{i}\}_{i=1}^{N}$ where each $\mathcal{G}_{i}=(V_{i},E_{i})$ consists of a set of nodes $V_{i}$ and a set of edges $E_{i}$ such that $n_{i}=|V_{i}|$ and $m_{i}=|E_{i}|$. We represent each $\mathcal{G}_{i}$ using its adjacency matrix ${\bf A}^{(i)}$. Further, each graph $\mathcal{G}_{i}$ has an associated target value, i.e., IPR value, $y^{(i)}\in\mathbb{R}$. For each node $v\in V_{i}$ in $\mathcal{G}_{i}$, there is an associated feature vector $\bm{h}_{j}^{(i)}\in\mathbb{R}^{d}$ and $\mathbf{H}^{(i)}\in\mathbb{R}^{|V_{i}|\times d}$ be the node feature matrix where $\bm{h}_{j}^{(i)}$ is the $j^{th}$ row. The objective is to learn a function $f:\mathcal{G}\to\mathbb{R}$, such that $f(\mathcal{G}_{i})\approx y^{(i)}$ for the given set of $N$ graphs.
@@ III Methodology and Results
The function $f$ can be parameterized by a model, in our case, Graph Convolutional Networks (GCN) and Graph Attention Networks (GAT) (Appendix B, C). Let $\bm{\theta}$ be the parameters of the model. The prediction for $\mathcal{G}_{i}$ is denoted by $\hat{y}^{(i)}=f(\mathcal{G}_{i};\bm{\theta})$. The model parameters $\bm{\theta}$ are learned by minimizing a loss function that measures the difference between the predicted values $\hat{y}^{(i)}$ and the true target values $y^{(i)}$. We use the Mean Squared Error (MSE) loss function as
@@
Hence, the graph regression problem can be formalized as finding the optimal parameters $\bm{\theta}$ of a model $f$ that minimize the loss function $\mathcal{L}(\bm{\theta})$, enabling the prediction of IPR values for given graphs.
@@ III.1 GCN Architecture
Figure 5: We use scale-free networks for training and testing. (a) We can observe very low accuracy during the test time for the GCN model. (b) However, we can observe increased accuracy using the Graph Attention Network. (c) We consider ER random and scale-free networks and associated IPR values leading to the dynamical states being delocalized and weakly localized. We can observe the GAT model predicts the state’s IPR value with significant accuracy. Here, we train the model for $500$ epochs.
@@
where $\bm{h}^{(3)}_{j}\in\mathbb{R}^{1\times k_{2}}$ is the $j^{th}$ row of ${\bf H}^{(3)}\in\mathbb{R}^{n\times k_{2}}$ and $\bm{z}\in\mathbb{R}^{k_{2}}$. Finally, we pass it to a linear part of the basic neural network (Fig. 3(e)) to perform regression task over graphs and output the predicted IPR value as
@@ III.2 Data Sets Preparation
For the regression task, we create the graph data sets by combining delocalized, weakly localized, and strongly localized network structures, which include star, wheel, path, cycle, and random graph models as Erdős-Rényi (ER) and the Scale-free (SF) networks (Appendix D). As predictions of the GNN model are independent of network size, we vary the size of the networks during dataset creation and store them as edge lists (${\bf A}^{(i)}$). For each network, we calculate the IPR value from the principal eigenvector (PEV) and assign it as the target value ($y^{(i)}$) to $\mathcal{G}_{i}$ in the datasets. Since we do not have predefined node features for the networks, and the GCN framework requires node features as input [25], we initialize the feature vector ($\bm{h}_{j}^{(i)}$) for each node with network-specific properties (clustering coefficient, PageRank, degree centrality, betweenness centrality and closeness centrality) to form the initial feature matrix (${\bf H}^{(0,i)}$) for the $i$th graph. However, the feature matrix could also be initialized with random binary values (i.e., 00, $1$) [25]. We pass the edge indices, feature matrices, and labels into the model for the regression task, $\{({\bf A}^{(i)},{\bf H}^{(0,i)},y^{(i)})\}_{i=1}^{N}$. We primarily use small-sized networks for training the model, while testing is conducted on networks ranging from small to large sizes, including sizes similar to and beyond those used in training.
[…]
We access the datasets through PyTorch Geometric libraries and preprocess the data sets by removing all the disconnected graphs and those with fewer than $10$ nodes. We incorporate node features extracted from network properties. The sizes of our preprocessed datasets are $\texttt{N}_{\texttt{ENZ}}=441$, $\texttt{N}_{\texttt{NCI}}=2796$, and $\texttt{N}_{\texttt{MCF}}=25084$ with each nodes having seven features as in the model network. Note that IPR values for disconnected graphs are trivially high, and we focus exclusively on connected graphs in our study. For real-world data sets, we divide $80\%$ of the data for training and the rest for testing.
@@ III.3 Training and Testing Strategy
During the training, the GCN model initializes the model parameters ($\bm{\theta}$). We initialize ${\bf W}^{(l-1)}$ at random using the initialization described in Glorot & Bengio (2010) [28]. During the forward pass, for each graph $\mathcal{G}_{i}$, we compute the graph representation using the GCN layers, which involves message passing and aggregation of node features (Fig. (3)). Finally, the model predicts the target value $\hat{y}^{(i)}$. Further, the model computes the loss $\mathcal{L}(\bm{\theta})$ using the predicted ($\hat{y}^{(i)}$) and true target ($y^{(i)}$) values. During the backward pass, the model computes the gradients of the loss with respect to the model parameters. In the next step, the model updates the parameters using an optimization algorithm such as Adam with a learning rate of $0.01$ and a weight decay of $45e-$ 4. We repeat the forward propagation, loss computation, backward propagation, and parameter update steps until convergence. For our experiment, we choose $d=7$ features for each node and weight matrix sizes as $k_{0}=k_{1}=k_{2}=64$ in different layers.
@@
We start by creating a simple experimental setup where the input dataset contains only two different types of model networks. One type of network (cycle) is associated with delocalized steady-state behavior, and another (star) is in the strongly localized behavior. During the training phase, we send the edge list (${\bf A}^{(i)}$), node feature matrix (${\bf H}^{(0,i)}$), and IPR values ($y^{(i)}$) associated with the graphs as labels for the regression task. Once the model is trained, one can observe that the GCN model accurately predicts the IPR value for the two different types of networks (Fig. 4(a)). More importantly, we train the model with smaller-size networks ($n_{i}=200$ to $300$) and test it with large-size networks ($n_{i}=400$ to $500$) and training datasets contains $N_{train}=1000$ networks and testing datasets size as $N_{test}=500$. Thus, the training cost would be less, and it can easily handle large networks. Furthermore, for the expressivity of the model, we increase the datasets by incorporating two more different types of graphs (path and wheel graphs), where one is delocalized and the other is in strongly localized structures, and we trained the model. We repeat the process by sending the datasets for the regression task to our model and observing that the model provides good accuracy for the test data sets (Fig. 4(b)). Finally, we apply the threshold function (Eq. 10) on the predicted values and achieve very high accuracy in identifying the dynamic state during the testing (Fig. 4(c, d)). One can observe that the GCN model learns the IPR value well for the above network structures.
@@
We move further and incorporate random graph structures (ER and scale-free random networks) in the data sets. Note that the ER random graph belongs to the delocalized state, and SF belongs to both the delocalized and weakly localized state. We train the model with only the SF networks, and during the testing time, one can observe accuracy is not good (Fig. 5(a)). To resolve this, we changed the model and the parameters. We choose the Graph Attention network [29], update the loss function by considering the $\log$ value, and choose AdamW optimizer instead of Adam. We also use a dropout rate of $0.6$ and set learning rare to $51e-$ 5 in the model. The new setup leads to improvement in the results (Fig. 5(b)). Now, we consider both ER and SF networks and train the model, and during the testing time, we can observe good accuracy in predicting the IPR values (Fig. 5(c)).

### v3

@@ I Introduction
Figure 1: Steady-state behavior in linear dynamical systems on complex networks. We depict the nodes in the graphs with xyx-y coordinates. We assign the sizes of a node based on the degree of a node. The zaxisz-axis portrays the amount of information ($x_{i}^{*}$) on a node in the steady state. The steady-state behavior of linear dynamics on the ER random network leads to delocalization, the scale-free network shows weak localization, and the star network shows a strong localization.
@@ II Problem Definition
Figure 2: The architecture of the Graph Neural Network (GNN) for the regression task over the graphs. The $i^{th}$ input graph (${\bf A}^{(i)}$) and the associated node features (${\bf H}^{(i)}$) are given in matrix form to the models. The GNN layers output an updated node feature matrix ${\bf H}^{{}^{\prime}(i)}$, which is passed through a readout layer to produce a graph-level representation $\bm{z}^{(i)}$. A fully connected layer then predicts the IPR value. Finally, we apply a threshold scheme (Eq. (10)) to identify different linear dynamical states.
We quantify the (de)localization behavior of the steady state ($\bm{x}^{*}$) or the PEV ($\bm{u}\equiv\bm{u}_{1}^{\mathbf{A}}$) using the inverse participation ratio (IPR), defined as the sum of the fourth powers of the state vector entries [20] as
@@
where $x^{*}_{i}$ is the $i^{\text{th}}$ component of $\bm{x^{*}}=(x^{*}_{1},x^{*}_{2},\dots,x^{*}_{n})^{T}$ and $\sum_{i=1}^{n}{x^{*}_{i}}^{2}$ is the normalization term. A vector with component $(c,c,\ldots,c)^{T}$ is delocalized and has $y_{\bm{x}^{*}}=\frac{1}{n}$ for some positive constant $c>0$, whereas the vector with components $(1,0,\ldots,0)^{T}$ yields $y_{\bm{x}^{*}}=1$ and referred as most localized. Furthermore, we consider the networks to be simple, connected and undirected [21]. Hence, some information can easily propagate from one node to another and we never get a steady-state vector of the form $\bm{x}^{*}=(1,0,\ldots,0)^{T}$ for a connected network, and thus the IPR value lies between $\frac{1}{n}\leq y_{\bm{x}^{*}}<1$. Therefore, the localization-delocalization behavior of the linear dynamics in the network is quantified using a real value, i.e., each graph associates an IPR value [22, 20, 8, 11]. We also employ entropy measurement for the localization, and more details are in Appendix F. To identify the states ($\bm{x}^{*}$) belong to which category of dynamical behavior for linear dynamics, we formalize a threshold scheme for identifying IPR values ($y\equiv y_{\bm{x}^{*}}$) lies in the range $[1/n,1)$. We define two thresholds, $\tau_{1}$ and $\tau_{2}$, such that $1/n\leq\tau_{1}<\tau_{2}<1$. An additional parameter $\epsilon$ ($\epsilon>0$) defines some flexibility around the thresholds.
@@
[…]
For instance, we consider a set of threshold values as $\tau_{1}=0.05$, $\tau_{2}=0.2$, $\epsilon=1e-6$. Now, if we consider a regular network (each node have the same degree) of $n$ nodes, we have PEV, $\bm{u}^{\mathcal{R}}=(\frac{1}{\sqrt{n}},\frac{1}{\sqrt{n}},\ldots,\frac{1}{\sqrt{n}})^{T}$ of A (Theorem 6 [21]) yielding, $y_{\bm{u}^{\mathcal{R}}}=\frac{1}{n}$, thus $y_{\bm{u}^{\mathcal{R}}}\rightarrow 0$ as $n\rightarrow\infty$. On the other hand, for a star graph having $n$ nodes, $\bm{u}^{\mathcal{S}}=\biggl{(}\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2(n-1)}},\ldots,\frac{1}{\sqrt{2(n-1)}}\biggr{)}^{T}$ and, $y_{\bm{u}^{\mathcal{S}}}=\frac{1}{4}+\frac{1}{4(n-1)}$. Hence, for $n\rightarrow\infty$, we get $y_{\bm{u}^{\mathcal{S}}}\approx 0.25$, and PEV is strongly localized for the star networks. Further, it is also difficult to find a closed functional form of PEV for any network, and thereby, it is hard to find the IPR value analytically. For instance, in Erdös-Rényi (ER) random networks, we get a delocalized PEV due to each node having the same expected degree [23]. In contrast, the presence of power-law degree distribution for SF networks leads to some localization in the PEV. For SF networks, the IPR value, while being larger than the ER random networks, is much lesser than the star networks [24]. It may seem that when the network structure is close to regular, linear dynamics are delocalized, and increasing degree heterogeneity increases the localization. However, always looking at the degree heterogeneity not able to decide localization and analyzing structural and spectral properties is essential [20, 11]. A fundamental question at the core of the structural-dynamical relation is: Can we predict the steady state behavior of a linear dynamical process in complex networks?
@@
Figure 4: We train the GCN model with two types of structures associated with delocalized and strongly localized states. The input to the GCN model is the cycle and star graphs and the associated target values, i.e., IPR. (a) We give datasets with varying-sized cycle and star networks during the test time. We can observe the true and predicted IPR values. Finally, we apply the threshold function (Eq. (10)) on the predicted IPR value. For our study, we choose the parameters for the threshold values as $\tau_{1}=0.05$, $\tau_{2}=0.2$, and $\epsilon=1e-6$ in Eq. (10). We mark the delocalized region ($r_{1}$) with an orange color box and the strongly localized region ($r_{3}$) with a red color box based on $\tau_{1}$ and $\tau_{2}$ values. This visualization enables us to identify the test network as delocalized and strongly localized based on their predicted IPR values. If predicted IPR values fall within the designated region, the identification of steady-state behavior prediction is correct. Notably, the threshold scheme also allows us to identify the correct behavior even when predicted IPR values deviate from the original values but lie within the threshold boundary. (b) To observe the expressivity of the model, we incorporate two other different network structures (wheel and path networks) and repeat the process. (c, d) We can observe the model-predicted values with high accuracy in the confusion matrix (in $\%$).
Here, we formulate the problem as a graph regression task to predict a target value, IPR, associated with each graph structure. For a given set of graphs $\{\mathcal{G}_{i}\}_{i=1}^{N}$ where each $\mathcal{G}_{i}=(V_{i},E_{i})$ consists of a set of nodes $V_{i}$ and a set of edges $E_{i}$ such that $n_{i}=|V_{i}|$ and $m_{i}=|E_{i}|$. We represent each $\mathcal{G}_{i}$ using its adjacency matrix ${\bf A}^{(i)}$. Further, each graph $\mathcal{G}_{i}$ has an associated target value, i.e., IPR value, $y^{(i)}\in\mathbb{R}$. For each node $v\in V_{i}$ in $\mathcal{G}_{i}$, there is an associated feature vector $\bm{h}_{j}^{(i)}\in\mathbb{R}^{1\times d}$ and $\mathbf{H}^{(i)}\in\mathbb{R}^{|V_{i}|\times d}$ be the node feature matrix where $\bm{h}_{j}^{(i)}$ is the $j^{th}$ row. The objective is to learn a function $f:\mathcal{G}\to\mathbb{R}$, such that $f(\mathcal{G}_{i})\approx y^{(i)}$ for the given set of $N$ graphs.
@@ III Methodology and Results
The function $f$ can be parameterized by a model, in our case, Graph Convolutional Networks (GCN) and Graph Attention Networks (GAT) (Appendices B, C). Let $\bm{\theta}$ be the parameters of the model. The prediction for $\mathcal{G}_{i}$ is denoted by $\hat{y}^{(i)}=f(\mathcal{G}_{i};\bm{\theta})$. The model parameters $\bm{\theta}$ are learned by minimizing a loss function that measures the difference between the predicted values $\hat{y}^{(i)}$ and the true target values $y^{(i)}$. We use the Mean Squared Error (MSE) loss function as
@@
To improve numerical stability, we apply a logarithmic transformation and compute the loss as
$\mathcal{L}=\text{MSE}(\log(1+\hat{y}),\log(1+y))$ (12)
This formulation helps compress the scale of the outputs and targets, particularly useful when the target values span multiple orders of magnitude. Hence, the graph regression problem can be formalized as finding the optimal parameters $\bm{\theta}$ of a model $f$ that minimize the loss function $\mathcal{L}(\bm{\theta})$, enabling the prediction of IPR values for given graphs.
@@ III.1 GCN Architecture
Figure 5: We use scale-free networks for training and testing. (a) We can observe very low accuracy during the test time for the GCN model. (b) However, we can observe increased accuracy using the Graph Attention Network. (c) We consider ER random and scale-free networks and associated IPR values leading to the dynamical states being delocalized and weakly localized. We can observe the GAT model predicts the state’s IPR value with significant accuracy. Here, we train the model for $200$ epochs.
@@
where $\bm{h}^{(3)}_{j}\in\mathbb{R}^{1\times k_{2}}$ is the $j^{th}$ row of ${\bf H}^{(3)}\in\mathbb{R}^{n\times k_{2}}$ and $\bm{z}\in\mathbb{R}^{1\times k_{2}}$. Finally, we pass it to a linear part of the basic neural network (Fig. 3(e)) to perform regression task over graphs and output the predicted IPR value as
@@ III.2 Data Sets Preparation
For the regression task, we create the graph data sets by combining delocalized, weakly localized, and strongly localized network structures, which include star, wheel, path, cycle, and random graph models as Erdős-Rényi (ER) and the scale-free (SF) networks (Appendix D). As predictions of the GNN model are independent of network size, we vary the size of the networks during dataset creation and store them as adjacency matrices (${\bf A}^{(i)}$). For each network, we calculate the IPR value from the principal eigenvector (PEV) and assign it as the target value ($y^{(i)}$) to $\mathcal{G}_{i}$ in the datasets. Since we do not have predefined node features for the networks, and the GCN framework requires node features as input [25], we initialize the feature vector ($\bm{h}_{j}^{(i)}$) for each node with network-specific properties (clustering coefficient, PageRank, degree centrality, betweenness centrality and closeness centrality) to form the initial feature matrix (${\bf H}^{(0,i)}$) for the $i$th graph. However, the feature matrix could also be initialized with random binary values (i.e., 0, 11) [25] (Appendix E). We pass the adjacency matrices, feature matrices, and labels into the model for the regression task, $\{({\bf A}^{(i)},{\bf H}^{(0,i)},y^{(i)})\}_{i=1}^{N}$. We primarily use small-sized networks for training the model, while testing is conducted on networks ranging from small to large sizes, including sizes similar to and beyond those used in training.
We also consider real-world benchmark datasets (ENZYMES, $\texttt{NCI}1$ and $\texttt{MCF}-7$) to train the model [26, 27]. ENZYMES is a dataset of $\texttt{N}_{\texttt{ENZ}}=600$ protein tertiary structures obtained from the BRENDA enzyme database. The $\texttt{NCI}1$ graph dataset is a benchmark dataset used in cheminformatics, where each graph represents a chemical compound, with nodes representing atoms and edges representing bonds between them. The $\texttt{NCI}1$ contains $\texttt{N}_{\texttt{NCI}}=4110$ graphs, and each node has $37$ features. The $\texttt{MCF}-7$ dataset consists of small molecule activities against breast cancer tumors of $\texttt{N}_{\texttt{MCF}}=27770$ graphs where each node has $46$ features. We access the datasets through PyTorch Geometric libraries and preprocess the data sets by removing all the disconnected graphs and those with fewer than $10$ nodes. We incorporate node features extracted from network-specific properties. The sizes of our preprocessed datasets are $\texttt{N}_{\texttt{ENZ}}=441$, $\texttt{N}_{\texttt{NCI}}=2796$, and $\texttt{N}_{\texttt{MCF}}=25084$ with each nodes having network-specific features as in the model network. Note that IPR values for disconnected graphs are trivially high, and we focus exclusively on connected graphs in our study. For real-world data sets, we divide $80\%$ of the data for training and the rest for testing.
@@ III.3 Training and Testing Strategy
During the training, the GCN model initializes the model parameters ($\bm{\theta}$). We initialize ${\bf W}^{(l-1)}$ at random using the initialization described in Glorot & Bengio (2010) [28]. During the forward pass, for each graph $\mathcal{G}_{i}$, we compute the graph representation using the GCN layers, which involves message passing and aggregation of node features (Fig. (3)). Finally, the model predicts the target value $\hat{y}^{(i)}$. Further, the model computes the loss $\mathcal{L}(\bm{\theta})$ using the predicted ($\hat{y}^{(i)}$) and true target ($y^{(i)}$) values. During the backward pass, the model computes the gradients of the loss with respect to the model parameters. In the next step, the model updates the parameters using an optimization algorithm such as Adam with a learning rate of $0.01$ and a weight decay of 5e45e-4. We repeat the forward propagation, loss computation, backward propagation, and parameter update steps until convergence. For our experiment, we choose weight matrix sizes as $k_{0}=k_{1}=k_{2}=64$ in different layers.
@@
We start by creating a simple experimental setup where the input dataset contains only two different types of model networks. One type of network (cycle) is associated with delocalized steady-state behavior, and another (star) is in the strongly localized behavior. During the training phase, we send the adjacency matrix (${\bf A}^{(i)}$), node feature matrix (${\bf H}^{(0,i)}$), and IPR values ($y^{(i)}$) associated with the graphs as labels for the regression task. Once the model is trained, one can observe that the GCN model accurately predicts the IPR value for the two different types of networks (Fig. 4(a)). More importantly, we train the model with smaller-size networks ($n_{i}=200$ to $300$) and test it with large-size networks ($n_{i}=400$ to $500$) and training datasets contains $N_{train}=1000$ networks and testing datasets size as $N_{test}=500$ where $50\%$ are cycle and $50\%$ are star networks. Thus, the training cost would be less, and it can easily handle large networks. Furthermore, for the expressivity of the model, we increase the datasets by incorporating two more different types of graphs (path and wheel graphs), where one is delocalized and the other is in strongly localized structures, and we trained the model. We repeat the process by sending the datasets for the regression task to our model and observing that the model provides good accuracy for the test data sets (Fig. 4(b)). Finally, we apply the threshold function (Eq. 10) on the predicted values and achieve very high accuracy in identifying the dynamic state during the testing (Fig. 4(c, d)). One can observe that the GCN model learns the IPR value well for the above network structures.
@@
We move further and incorporate random graph structures (ER and scale-free random networks) in the data sets. Note that the ER random graph belongs to the delocalized state, and SF belongs to both the delocalized and weakly localized state. We train the model with only the SF networks, and during the testing time, one can observe accuracy is not good (Fig. 5(a)). To resolve this, we changed the model and the parameters. We choose the Graph Attention network [29], update the loss function by considering the $\log$ value (Eq. 12), and choose AdamW optimizer instead of Adam. We also use a dropout rate of $0.6$ and set learning rare to 1e51e-5 in the model. The new setup leads to improvement in the results (Fig. 5(b)). Now, we consider both ER and SF networks and train the model, and during the testing time, we can observe good accuracy in predicting the IPR values (Fig. 5(c)). We also consider other graph models (Fig. A2(a, c)) and entropy measures for localization (Fig. A2(b, d)) and observe that our models perform well (Appendices D, F).
Figure 6: Scale invariant. Confusion matrices for the GAT model trained on medium-sized networks (size 5001500500-1500) and tested on significantly larger networks (size 5000150005000-15000). (a) Binary classification between delocalized and localized states achieves $95\%$ accuracy. (b) Three-class classification (delocalized ($r_{1}\mapsto 0$), weakly localized ($r_{2}\mapsto 1$), and strongly localized ($r_{3}\mapsto 2$)) achieves $96\%$ accuracy, demonstrating scale-invariant performance of the model.